Tom has 37 coins, consisting only of dimes and quarters, worth $7.45. How many dimes and how many quarters does he have?
Answer by checkley77(12844)

q+d=37
q=37-d
.25q+.10d=7.45
.25(37-d)+.10d=7.45
9.25-.25d+.10d=7.45
-.15d=7.45-9.25
-.15d=-1.80
d=-1.80/.15
d=12 dimes.
q=37-12
q=25 quarters.
Proof:
.25*25=.10*12=7.45
6.25+1.20=7.45
7.45=7.45